OpenStudy (anonymous):

2 and 2 over 7 ÷ 4 over 7

sammixboo (sammixboo):

\(\sf \large \color{magenta}{2\frac{2}{7}÷\frac{4}{7}}\) Does your problem look like this?

sammixboo (sammixboo):

\(\sf \large \color{magenta}{2\frac{2}{7}/\frac{4}{7}}\) (the / is a division symbol)